40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1021     Entry time: Mon Sep 6 16:37:05 2010
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: New gyro noise spectrum 

Now that stuff is working again, I took a spectrum of the gyro in the newest configuration (the one detailed here).

The calibration is: 6.103 x 10-4 V/ct * 375 kHz/V * (lambda * S)/(4*A) [rad/s / Hz] = 3.09 x 10-4 rad/s /ct

gyro_noise_9_3_2010.png

It is roughly a factor of 10 better than before at and below 10 Hz, though this comes with the grain of salt that the calibration was off by a factor of 2*pi before, and that there exists some ambiguity over whether or not the measurement can be trusted due to our incomplete understanding of the loop. The noise above 10 Hz is pretty much the same.

It appears that someone was futzing with the gyro over the week, as the BOOSTs on both boxes were turned OFF and the CW camera was off center. Please document your futzings.

ELOG V3.1.3-